OpenStudy (anonymous):

Suppose Galileo and his assistant were 1 km apart when Galileo uncovered his lantern to try to measure the speed of light. How long would it have taken the light signal to reach his assistant? A. 3.0 * 10^8 s B. 3.3 * 10^-9 s C. 1.0 s D. 3.3 * 10^-6 s

OpenStudy (michele_laino):

it is simple, since we have to apply this formula: \[t = \frac{d}{c} = \frac{{10 \times 3}}{{3 \times {{10}^8}}}\]

OpenStudy (michele_laino):

oops..I have made a typo, here is the right formula:

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

\[t = \frac{d}{c} = \frac{{{{10}^3}}}{{3 \times {{10}^8}}}\]

OpenStudy (anonymous):

we get 3.333333333E-6?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

so our solution is choice D?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay!! :) Thanks!!

OpenStudy (michele_laino):

:)